OpenStudy (anonymous):

A car is traveling at 60.6 mph on a horizontal highway. The acceleration of gravity is 9.8 m/s2. If the coefficient of friction between road and tires on a rainy day is 0.056, what is the minimum distance in which the car will stop? Answer in units of m. m = 60-kg.

OpenStudy (anonymous):

for the min distance condition : you must be knowing that maximum/limiting friction is acting. It'll provide a deceleration of (coeff)*g . Please convey any doubts.

OpenStudy (anonymous):

you want direct answer or procedure

OpenStudy (anonymous):

IF the friction coefficient stays constant during braking then: 60,6mph=27m/s B - friction coefficient F(friction)=B*N, where N is the normal reaction, I assume that the car is on horisontal road, so N=m*g, where m is the mass of the car. F(braking)=m*a, where a is the acceleration. On a limiting case, where F(braking)=F(friction), we get a=B*g. From the equation x=v^2/2*a (because v(final)=0), we get x=(v^2)/(2*B*g)=660m In conclusion, this friction coefficient is ridiculously low, hence the long distance to stop.

OpenStudy (anonymous):

what is B?

OpenStudy (anonymous):

he meant B = friction coefficient

OpenStudy (anonymous):

Ok thank you @kappa007

OpenStudy (anonymous):

:)